A microprocessor uses RAM chips of 1×1024 capacity.

How many chips will be required to obtain a memory capacity of 16 𝐾 bytes?

How many chips will be required and how many address lines will be connected to provide a capacity of 1024 bytes?


1
Expert's answer
2021-01-01T13:37:08-0500

We are given:

The capacity of the chip is "= (1024 \times 1 ) = 1024" bits

We have to obtain a memory whose capacity is "16" K bytes or "(16 \times 1024 \times 8 )" bits.

Hence:

The number of chips required to obtain a memory capacity of "16" K bytes are:

"= \dfrac{16 \times 1024 \times 8 }{1024}"

"\, = \color{blue}{128}"

Again:

We have to obtain a memory whose capacity is "1024" bytes or "(1024 \times 8)" bits

Hence:

The number of chips required to obtain a memory capacity of "1024" bytes are:

"= \dfrac{1024 \times 8}{1024}"

"= \color{blue}{ 8}"

Now:

Since "2^{10} = 1024" so a memory chip capacity of "1024" bytes connected with "\color{blue}{10}" address lines.